The Linux XFree86 HOWTO

by Matt Welsh, mdw@sunsite.unc.edu v3.0, 15 March 1995

This document describes how to obtain, install, and configure version 3.1.1 of the XFree86 version of the X Window System (X11R6) for Linux systems. It is a step-by-step guide to configuring XFree86 on your system.

Contents

1 Introduction 1

2 Hardware requirements 2

3 Installing XFree86 4

4 Configuring XFree86 7

5 Filling in video card information 14

6 Running XFree86 17

7 Running Into Trouble 18

8 Copyright 19

1 Introduction

The X Window System is a large and powerful (and somewhat complex) graphics environment for UNIX systems. The original X Window System code was developed at MIT; commercial vendors have since made X the industry standard for UNIX platforms. Virtually every UNIX workstation in the world runs some variant of the X Window system.

A free port of the MIT X Window System version 11, release 6 (X11R6) for 80386/80486/Pentium UNIX systems has been developed by a team of programmers originally headed by David Wexelblat (dwex@XFree86.org). The release, known as XFree86, is available for System V/386, 386BSD, and other x86 UNIX implementations, including Linux. It includes all of the required binaries, support files, libraries, and tools.

In this document, we’ll give a step-by-step description of how to install and configure XFree86 for Linux, but you will have to fill in some of the details yourself by reading the documentation released with XFree86 itself. (This documentation is discussed below.) However, using and customizing the X Window System is far beyond the scope of this document—for this purpose you should obtain one of the many good books on using the X Window System.

2 Hardware requirements

As of XFree86 version 3.1.1, released in February 1995, the following video chipsets are supported. The documentation included with your video adaptor should specify the chipset used. If you are in the market for a new video card, or are buying a new machine that comes with a video card, have the vendor find out exactly what the make, model, and chipset of the video card is. This may require the vendor to call technical support on your behalf; in general vendors will be happy to do this. Many PC hardware vendors will state that the video card is a “standard SVGA card” which “should work” on your system. Explain that your software (mention Linux and XFree86!) does not support all video chipsets and that you must have detailed information. You can also determine your videocard chipset by running the SuperProbe program included with the XFree86 distribution. This is covered in more detail below. The following standard SVGA chipsets are supported:

• Tseng ET3000, ET4000AX, ET4000/W32

• Western Digital/Paradise PVGA1

• Western Digital WD90C00, WD90C10, WD90C11, WD90C24, WD90C30, WD90C31, WD90C33 • Genoa GVGA

• Trident TVGA8800CS, TVGA8900B, TVGA8900C, TVGA8900CL, TVGA9000, TVGA9000i, TVGA9100B, TVGA9200CX, TVGA9320, TVGA9400CX, TVGA9420

• ATI 18800, 18800-1, 28800-2, 28800-4, 28800-5, 28800-6, 68800-3, 68800-6, 68800AX, 68800LX, 88800 • NCR 77C22, 77C22E, 77C22E+

• Cirrus Logic CLGD5420, CLGD5422, CLGD5424, CLGD5426, CLGD5428, CLGD5429, CLGD5430, CLGD5434, CLGD6205, CLGD6215, CLGD6225, CLGD6235, CLGD6420

• Compaq AVGA

• OAK OTI067, OTI077

• Avance Logic AL2101

• MX MX68000, MX680010

• Video 7/Headland Technologies HT216-32 The following SVGA chipsets with accelerated features are also supported:

• 8514/A (and true clones)

• ATI Mach8, Mach32

• Cirrus CLGD5420, CLGD5422, CLGD5424, CLGD5426, CLGD5428, CLGD5429, CLGD5430, CLGD5434, CLGD6205, CLGD6215, CLGD6225, CLGD6235

• S3 86C911, 86C924, 86C801, 86C805, 86C805i, 86C928, 86C864, 86C964

• Western Digital WD90C31, WD90C33

• Weitek P9000

• IIT AGX-014, AGX-015, AGX-016

• Tseng ET4000/W32, ET4000/W32i, ET4000/W32p

Video cards using these chipsets are supported on all bus types, including VLB and PCI.

All of the above are supported in both 256 color and monochrome modes, with the exception of the Avance Logic, MX and Video 7 chipsets, which are only supported in 256 color mode. If your video card has enough DRAM installed, many of the above chipsets are supported in 16 and 32 bits-per-pixel mode (specifically, some Mach32, P9000, S3 and Cirrus boards). The usual configuration is 8 bits per pixel (that is, 256 colors).

The monochrome server also supports generic VGA cards, the Hercules monochrome card, the Hyundai HGC1280, Sigma LaserView, and Apollo monochrome cards. On the Compaq AVGA, only 64k of video memory is supported for the monochrome server, and the GVGA has not been tested with more than 64k.

This list will undoubtedly expand as time passes. The release notes for the current version of XFree86 should contain the complete list of supported video chipsets.

One problem faced by the XFree86 developers is that some video card manufacturers use non-standard mechanisms for determining clock frequencies used to drive the card. Some of these manufacturers either don’t release specifications describing how to program the card, or they require developers to sign a nondisclosure statement to obtain the information. This would obviously restrict the free distribution of the XFree86 software, something that the XFree86 development team is not willing to do. For a long time, this has been a problem with certain video cards manufactured by Diamond, but as of release 3.1 of XFree86, Diamond has started to work with the development team to release free drivers for these cards.

The suggested setup for XFree86 under Linux is a 486 machine with at least 8 megabytes of RAM, and a video card with a chipset listed above. For optimal performance, we suggest using an accelerated card, such as an S3-chipset card. You should check the documentation for XFree86 and verify that your particular card is supported before taking the plunge and purchasing expensive hardware. Benchmark ratings comparisons for various video cards under XFree86 are posted routinely to the USENET newsgroups

comp.windows.x.i386unix and comp.os.linux.x.

As a side note, my personal Linux system is a 486DX2-66, 20 megabytes of RAM, and is equipped with a VLB S3-864 chipset card with 2 megabytes of DRAM. I have run X benchmarks on this machine as well as on Sun Sparc IPX workstations. The Linux system is roughly 7 times faster than the Sparc IPX (for the curious, XFree86-3.1 under Linux, with this video card, runs at around 171,000 xstones; the Sparc IPX at around 24,000). In general, XFree86 on a Linux system with an accelerated SVGA card will give you much greater performance than that found on commercial UNIX workstations (which usually employ simple framebuffers for graphics).

Your machine will need at least 4 megabytes of physical RAM, and 16 megabytes of virtual RAM (for example, 8 megs physical and 8 megs swap). Remember that the more physical RAM that you have, the less that the system will swap to and from disk when memory is low. Because swapping is inherently slow (disks are very slow compared to memory), having 8 megabytes of RAM or more is necessary to run XFree86 comfortably. A system with 4 megabytes of physical RAM could run much (up to 10 times) more slowly than one with 8 megs or more.

3 Installing XFree86

The Linux binary distribution of XFree86 can be found on a number of FTP sites. On sunsite.unc.edu, it is found in the directory /pub/Linux/X11. (As of the time of this writing, the current version is 3.1.1; newer versions are released periodically). It’s quite likely that you obtained XFree86 as part of a Linux distribution, in which case downloading the software separately is not necessary. If you are downloading XFree86 directly, This table lists the files in the XFree86-3.1 distribution. One of the following servers is required:

XF86-3.1.1-8514.tar.gz

Server for 8514-based boards. XF86-3.1.1-AGX.tar.gz

Server for AGX-based boards. XF86-3.1.1-Mach32.tar.gz

Server for Mach32-based boards. XF86-3.1.1-Mach8.tar.gz

Server for Mach8-based boards. XF86-3.1.1-Mono.tar.gz

Server for monochrome video modes. XF86-3.1.1-P9000.tar.gz

Server for P9000-based boards. XF86-3.1.1-S3.tar.gz

Server for S3-based boards. XF86-3.1.1-SVGA.tar.gz

Server for Super VGA-based boards. XF86-3.1.1-VGA16.tar.gz

Server for VGA/EGA-based boards.

XF86-3.1.1-W32.tar.gz Server for ET4000/W32-based boards.

All of the following files are required:

XF86-3.1.1-bin.tar.gz

The rest of the X11R6 binaries. XF86-3.1.1-cfg.tar.gz

Config files for xdm, xinit and fs. XF86-3.1.1-doc.tar.gz

Documentation and manpages. XF86-3.1.1-inc.tar.gz

Include files. XF86-3.1.1-lib.tar.gz

Shared X libraries and support files. XF86-3.1-fnt.tar.gz

Basic fonts.

The following files are optional:

XF86-3.1-ctrb.tar.gz

Selected contrib programs. XF86-3.1-extra.tar.gz

Extra XFree86 servers and binaries. XF86-3.1-lkit.tar.gz

Server linkkit for customization. XF86-3.1-fnt75.tar.gz

75-dpi screen fonts. XF86-3.1-fnt100.tar.gz

100-dpi screen fonts. XF86-3.1-fntbig.tar.gz

Large Kanji and other fonts. XF86-3.1-fntscl.tar.gz

Scaled fonts (Speedo, Type1).

XF86-3.1-man.tar.gz Manual pages.

XF86-3.1-pex.tar.gz PEX binaries, includes and libraries.

XF86-3.1-slib.tar.gz Static X libraries and support files.

XF86-3.1-usrbin.tar.gz Daemons which reside in /usr/bin.

XF86-3.1-xdmshdw.tar.gz Shadow password version of xdm.

The XFree86 directory should contain README files and installation notes for the current version.

All that is required to install XFree86 is to obtain the above files, create the directory /usr/X11R6 (as root), and unpack the files from /usr/X11R6 with a command such as:

     gzip -dc XF86-3.1.1-bin.tar.gz | tar xfB -

Remember that these tar files are packed relative to /usr/X11R6. so it’s important to unpack the files there.

After unpacking the files, you first need to link the file /usr/X11R6/bin/X to the server that you’re using. For example, if you wish to use the SVGA color server, /usr/bin/X11/X should be linked to /usr/X11R6/bin/XF86 SVGA. If you wish to use the monochrome server instead, relink this file to XF86 MONO with the command

     ln -sf /usr/X11R6/bin/XF86_MONO /usr/X11R6/bin/X

The same holds true if you are using one of the other servers.

If you aren’t sure which server to use, or don’t know your video card chipset, you can run the SuperProbe program found in /usr/X11R6/bin (included in the XF86-3.1-bin listed above). This program will attempt to determine your video chipset type and other information; write down its output for later reference.

You need to make sure that /usr/X11R6/bin is on your path. This can be done by editing your system default /etc/profile or /etc/csh.login (based on the shell that you, or other users on your system, use). Or you can simply add the directory to your personal path by modifying /etc/.bashrc or /etc/.cshrc, based on your shell.

You also need to make sure that /usr/X11R6/lib can be located by ld.so, the runtime linker. To do this, add the line

     /usr/X11R6/lib
to the file /etc/ld.so.conf, and run /sbin/ldconfig, as root.

4 Configuring XFree86

Setting up XFree86 is not difficult in most cases. However, if you happen to be using hardware for which drivers ar under development, or wish to obtain the best performance or resolution from an accelerated graphics card, configuring XFree86 can be somewhat time-consuming.

In this section we will describe how to create and edit the XF86Config file, which configures the XFree86 server. In many cases it is best to start out with a “basic” XFree86 configuration, one which uses a low resolution, such as 640x480, which should be supported on all video cards and monitor types. Once you have XFree86 working at a lower, standard resolution, you can tweak the configuration to exploit the capabilities of your video hardware. The idea is that you want to know that XFree86 works at all on your system, and that something isn’t wrong with your installation, before attempting the sometimes difficult task of setting up XFree86 for real use.

In addition to the information listed here, you should read the following documentation:

• The XFree86 documentation in /usr/X11R6/lib/X11/doc (contained within the XFree86-3.1-doc package). You should especially see the file README.Config, which is an XFree86 configuration tutorial.

• Several video chipsets have separate README files in the above directory (such as README.Cirrus and README.S3). Read one of these if applicable.

• The man page for XFree86.

• The man page for XF86Config.

• The man page for the particular server that you are using (such as XF86 SVGA or XF86 S3).

The main XFree86 configuration file is /usr/X11R6/lib/X11/XF86Config. This file contains information on your mouse, video card parameters, and so on. The file XF86Config.eg is provided with the XFree86 distribution as an example. Copy this file to XF86Config and edit it as a starting point.

The XF86Config man page explains the format of this file in detail. Read this man page now, if you have not done so already.

We are going to present a sample XF86Config file, piece by piece. This file may not look exactly like the sample file included in the XFree86 distribution, but the structure is the same.

Note that the XF86Config file format may change with each version of XFree86; this information is only valid for XFree86 version 3.1.

Also, you should not simply copy the configuration file listed here to your own system and attempt to use it. Attempting to use a configuration file which doesn’t correspond to your hardware could drive the monitor at a frequency which is too high for it; there have been reports of monitors (especially fixed-frequency monitors) being damaged or destroyed by using an incorrectly configured XF86Config file. The bottom line is this: Make absolutely sure that your XF86Config file corresponds to your hardware before you attempt to use it.

Each section of the XF86Config file is surrounded by the pair of lines

     Section "section-name"
       ...
     EndSection

The first part of the XF86Config file is Files, which looks like this:

     Section "Files"
         RgbPath     "/usr/X11R6/lib/X11/rgb"
         FontPath    "/usr/X11R6/lib/X11/fonts/misc/"
         FontPath    "/usr/X11R6/lib/X11/fonts/75dpi/"
     EndSection

The RgbPath line sets the path to the X11R6 RGB color database, and each FontPath line sets the path to a directory containing X11 fonts. In general you shouldn’t have to modify these lines; just be sure that there is a FontPath entry for each font type that you have installed (that is, for each directory in

/usr/X11R6/lib/X11/fonts).

The next section is ServerFlags, which specifies several global flags for the server. In general this section is empty.

     Section "ServerFlags"
     # Uncomment this to cause a core dump at the spot where a signal is
     # received.  This may leave the console in an unusable state, but may
     # provide a better stack trace in the core dump to aid in debugging
     #    NoTrapSignals
     # Uncomment this to disable the <Crtl><Alt><BS> server abort sequence
     #    DontZap
     EndSection

Here, we have all lines within the section commented out.

The next section is Keyboard. This should be fairly intuitive.

     Section "Keyboard"
         Protocol    "Standard"
         AutoRepeat  500 5
         ServerNumLock
     EndSection

Other options are available as well—see the XF86Config file if you wish to modify the keyboard configuration. The above should work for most systems.

The next section is Pointer which specifies parameters for the mouse device.

     Section "Pointer"
         Protocol    "MouseSystems"
         Device      "/dev/mouse"
     # Baudrate and SampleRate are only for some Logitech mice
     #    BaudRate   9600
     #    SampleRate 150

     # Emulate3Buttons is an option for 2-button Microsoft mice
     #    Emulate3Buttons
     # ChordMiddle is an option for some 3-button Logitech mice
     #    ChordMiddle
     EndSection

The only options that you should concern yourself with now are Protocol and Device. Protocol specifies the mouse protocol that your mouse uses (not the make or brand of mouse). Valid types for Protocol (under Linux—there are other options available for other operating systems) are:

   • BusMouse
   • Logitech
   • Microsoft
   • MMSeries
   • Mouseman
   • MouseSystems
   • PS/2
   • MMHitTab

BusMouse should be used for the Logitech busmouse. Note that older Logitech mice should use Logitech, but newer Logitech mice use either Microsoft or Mouseman protocols. This is a case in which the protocol doesn’t necessarily have anything to do with the make of the mouse.

Device specifies the device file where the mouse can be accessed. On most Linux systems, this is /dev/mouse. /dev/mouse is usually a link to the appropriate serial port (such as /dev/cua0) for serial mice, or to the appropriate busmouse device for busmice. At any rate, be sure that the device file listed in Device exists.

The next section is Monitor, which specifies the characteristics of your monitor. As with other sections in the XF86Config file, there may be more than one Monitor section. This is useful if you have multiple monitors connected to a system, or use the same XF86Config file under multiple hardware configurations. In general, though, you will need a single Monitor section.

     Section "Monitor"
         Identifier  "CTX 5468 NI"
         # These values are for a CTX 5468NI only! Don't attempt to use
         # them with your monitor (unless you have this model)
         Bandwidth    60

         HorizSync    30-38,47-50
         VertRefresh  50-90
         # Modes: Name      dotclock  horiz                vert
         ModeLine "640x480"  25       640 664 760 800      480 491 493 525
         ModeLine "800x600"  36       800 824 896 1024     600 601 603 625
         ModeLine "1024x768" 65       1024 1088 1200 1328  768 783 789 818
     EndSection

The Identifier line is used to give an arbitrary name to the Monitor entry. This can be any string; you will use it to refer to the Monitor entry later in the XF86Config file.

they are listed below.

HorizSync specifies the valid horizontal sync frequencies for your monitor, in kHz. If you have a multisync monitor, this can be a range of values (or several comma-separated ranges), as seen above. If you have a fixed-frequency monitor, this will be a list of discrete values, such as:

         HorizSync    31.5, 35.2, 37.9, 35.5, 48.95

Your monitor manual should list these values in the technical specifications section. If you do not have this information available, you should either contact the manufacturer or vendor of your monitor to obtain it. There are other sources of information, as well;

VertRefresh specifies the valid vertical refresh rates (or vertical synchronization frequencies) for your monitor, in Hz. Like HorizSync this can be a range or a list of discrete values; your monitor manual should list them.

HorizSync and VertRefresh are used only to double-check that the monitor resolutions that you specify are in valid ranges. This is to reduce the chance that you will damage your monitor by attempting to drive it at a frequency for which it was not designed.

The ModeLine directive is used to specify a single resolution mode for your monitor. The format of ModeLine is

     ModeLine name clock horiz-values vert-values

name is an arbitrary string, which you will use to refer to the resolution mode later in the file. dot-clock is the driving clock frequency, or “dot clock” associated with the resolution mode. A dot clock is usually specified in MHz, and is the rate at which the video card must send pixels to the monitor at this resolution. horiz-values and vert-values are four numbers each which specify when the electron gun of the monitor should fire, and when the horizontal and vertical sync pulses fire during a sweep.

How can you determine the ModeLine values for your monitor? The file VideoModes.doc, included with the XFree86 distribution, describes in detail how to determine these values for each resolution mode that your monitor supports. First of all, clock must correspond to one of the dot clock values that your video card can produce. Later in the XF86Config file you will specify these clocks; you can only use video modes which have a clock value supported by your video card.

There are two files included in the XFree86 distribution which may include ModeLine data for your monitor.

These files are modeDB.txt and Monitors, both of which are found in /usr/X11R6/lib/X11/doc.

You should start with ModeLine values for the VESA standard monitor timings, which most monitors support. modeDB.txt includes timing values for VESA standard resolutions. In that file, you will see entries such as

     # 640x480@60Hz Non-Interlaced mode
     # Horizontal Sync = 31.5kHz
     # Timing: H=(0.95us, 3.81us, 1.59us), V=(0.35ms, 0.064ms, 1.02ms)
     #
     # name        clock   horizontal timing     vertical timing      flags
      "640x480"     25.175  640  664  760  800    480  491  493  525

This is a VESA standard timing for a 640x480 video mode. It uses a dot clock of 25.175, which your video card must support to use this mode (more on this later). To include this entry in the XF86Config file, you’d use the line

     ModeLine "640x480" 25.175 640 664 760 800 480 491 493 525

Note that the name argument to ModeLine (in this case "640x480") is an arbitrary string—the convention is to name the mode after the resolution, but name can technically be anything descriptive which describes the mode to you.

For each ModeLine used the server will check that the specifications for the mode fall within the range of values specified with Bandwidth, HorizSync and VertRefresh. If they do not, the server will complain when you attempt to start up X (more on this later). For one thing, the dot clock used by the mode should not be greater than the value used for Bandwidth. (However, in many cases it is safe to use modes with a slightly higher bandwidth than your monitor can support.)

If the VESA standard timings do not work for you (you’ll know after trying to use them later) then the files modeDB.txt and Monitors include specific mode values for many monitor types. You can create ModeLine entries from the values found in those two files as well. Be sure to only use values for the specific model of monitor that you have. Note that many 14 and 15-inch monitors cannot support higher resolution modes, and often resolutions of 1024x768 at low dot clocks. This means that if you can’t find high resolution modes for your monitor in these files, then your monitor probably does not support those resolution modes.

If you are completely at a loss, and can’t find working ModeLine values for your monitor, you can follow the instructions in the VideoModes.doc file included in the XFree86 distribution to generate ModeLine values from the specifications listed in your monitor’s manual. While your mileage will certainly vary when attempting to generate ModeLine values by hand, this is a good place to look if you can’t find the values that you need. VideoModes.doc also describes the format of the ModeLine directive and other aspects of the XFree86 server in gory detail.

Lastly, if you do obtain ModeLine values which are almost, but not quite, right, then it may be possible to simply modify the values slightly to obtain the desired result. For example, if while running XFree86 the image on the monitor is shifted slightly, or seems to “roll”, you can follow the instructions in the VideoModes.doc file to try to fix these values. Also, be sure to check the knobs and controls on the monitor itself! In many cases it is necessary to change the horizontal or vertical size of the display after starting up XFree86 in order for the image to be centered and be of the appropriate size. Having these controls on the front of the monitor can certainly make life easier.

You shouldn’t use monitor timing values or ModeLine values for monitors other than the model that you own. If you attempt to drive the monitor at a frequency for which it was not designed, you can damage or even destroy it.

The next section of the XF86Config file is Device, which specifies parameters for your video card. Here is an example.

     Section "Device"
             Identifier "#9 GXE 64"
             # Nothing yet; we fill in these values later.
     EndSection

This section defines properties for a particular video card. Identifier is an arbitrary string describing the card; you will use this string to refer to the card later.

Initially, you don’t need to include anything in the Device section, except for Identifier. This is because we will be using the X server itself to probe for the properties of the video card, and entering them into the Device section later. The XFree86 server is capable of probing for the video chipset, clocks, RAMDAC, and amount of video RAM on the board.

Before we do this, however, we need to finish writing the XF86Config file. The next section is Screen, which specifies the monitor/video card combination to use for a particular server.

      Section "Screen"
          Driver     "Accel"
          Device     "#9 GXE 64"
          Monitor    "CTX 5468 NI"
          Subsection "Display"
              Depth      16
              Modes      "1024x768" "800x600" "640x480"
              ViewPort   0 0
              Virtual    1024 768
          EndSubsection
      EndSection

The Driver line specifies the X server that you will be using. The value values for Driver are:

   • Accel: For the XF86 S3, XF86 Mach32, XF86 Mach8, XF86 8514, XF86 P9000, XF86 AGX, and XF86 W32

servers;

• SVGA: For the XF86 SVGA server;

• VGA16: For the XF86 VGA16 server; • VGA2: For the XF86 Mono server;

• Mono: For the non-VGA monochrome drivers in the XF86 Mono and XF86 VGA16 servers.

You should be sure that /usr/X11R6/bin/X is a symbolic link to the server that you are using.

The Device line specifies the Identifier of the Device section corresponding to the video card to use for this server. Above, we created a Device section with the line

Identifier "#9 GXE 64"

Therefore, we use "#9 GXE 64" on the Device line here.

Similarly, the Monitor line specifies the name of the Monitor section to be used with this server. Here, "CTX 5468 NI" is the Identifier used in the Monitor section described above.

Subsection "Display" defines several properties of the XFree86 server corresponding to your monitor/video card combination. The XF86Config file describes all of these options in detail; most of them are icing on the cake and not necessary to get the system working.

The options that you should know about are:

• Depth. Defines the number of color planes—the number of bits per pixel. Usually, Depth is set to 8. For the VGA16 server, you would use a depth of 4, and for the monochrome server a depth of 1. If you are using an accelerated video card with enough memory to support more bits per pixel, you can set Depth to 16, 24, or 32. If you have problems with depths higher than 8, set it back to 8 and attempt to debug the problem later.

• Modes. This is the list of video mode names which have been defined using the ModeLine directive in the Monitor section. In the above section, we used ModeLines named "1024x768", "800x600", and "640x480". Therefore, we use a Modes line of

              Modes    "1024x768" "800x600" "640x480"

The first mode listed on this line will be the default when XFree86 starts up. After XFree86 is running, you can switch between the modes listed here using the keys ctrl-alt-numeric + and

ctrl-alt-numeric -.

It might be best, when initially configuring XFree86, to use lower resolution video modes, such as 640x480, which tend to work on most systems. Once you have the basic configuration working you can modify XF86Config to support higher resolutions.

• Virtual. Sets the virtual desktop size. XFree86 has the ability to use any additional memory on your video card to extend the size of your desktop. When you move the mouse pointer to the edge of the display, the desktop will scroll, bringing the additional space into view. Therefore, even if you are running at a lower video resolution such as 800x600, you can set Virtual to the total resolution which your video card can support (a 1-megabyte video card can support 1024x768 at a depth of 8 bits per pixel; a 2-megabyte card 1280x1024 at depth 8, or 1024x768 at depth 16). Of course, the entire area will not be visible at once, but it can still be used.

The Virtual feature is a nice way to utilize the memory of your video card, but it is rather limited. If you want to use a true virtual desktop, we suggest using fvwm, or a similar window manager, instead. fvwm allows you to have rather large virtual desktops (implemented by hiding windows, and so forth, instead of actually storing the entire desktop in video memory at once). See the man pages for fvwm for more details about this; most Linux systems use fvwm by default.

• ViewPort. If you are using the Virtual option described above, ViewPort sets the coordinates of the upper-left-hand corner of the virtual desktop when XFree86 starts up. Virtual 0 0 is often used; if this is unspecified then the desktop is centered on the virtual desktop display (which may be undesirable to you).

Many other options for this section exist; see the XF86Config man page for a complete description. In practice these other options are not necessary to get XFree86 initially working.

5 Filling in video card information

Your XF86Config file is now ready to go, with the exception of complete information on the video card. What we’re going to do is use the X server to probe for the rest of this information, and fill it into XF86Config.

Instead of probing for this information with the X server, the XF86Config values for many cards are listed in

the files modeDB.txt, AccelCards, and Devices. These files are all found in /usr/X11R6/lib/X11/doc. In

addition, there are various README files for certain chipsets. You should look in these files for information on your video card, and use that information (the clock values, chipset type, and any options) in the XF86Config file. If any information is missing, you can probe for it as described here.

In these examples we will demonstrate configuration for a #9 GXE 64 video card, which uses the XF86 S3 chipset. This card happens to be the one which the author uses, but the discussion here applies to any video card.

The first thing to do is to determine the video chipset used on the card. Running SuperProbe (found in /usr/X11R6/bin) will tell you this information, but you need to know the chipset name as it is known to the X server.

To do this, run the command

     X -showconfig

This will give the chipset names known to your X server. (The man pages for each X server list these as well.) For example, with the accelerated XF86 S3 server, we obtain:

     XFree86 Version 3.1 / X Window System
     (protocol Version 11, revision 0, vendor release 6000)
     Operating System: Linux
     Configured drivers:
       S3: accelerated server for S3 graphics adaptors (Patchlevel 0)
           mmio_928, s3_generic

The valid chipset names for this server are mmio 928 and s3 generic. The XF86 S3 man page describes these chipsets and which videocards use them. In the case of the #9 GXE 64 video card, mmio 928 is appropriate.

If you don’t know which chipset to use, the X server can probe it for you. To do this, run the command

     X -probeonly > /tmp/x.out 2>&1

if you use bash as your shell. If you use csh, try:

     X -probeonly &> /tmp/x.out

You should run this command while the system is unloaded, that is, while no other activity is occurring on the system. This command will also probe for your video card dot clocks (as seen below), and system load can throw off this calculation.

The output from the above (in /tmp/x.out should contain lines such as the following:

     XFree86 Version 3.1 / X Window System
     (protocol Version 11, revision 0, vendor release 6000)
     Operating System: Linux
     Configured drivers:
       S3: accelerated server for S3 graphics adaptors (Patchlevel 0)
           mmio_928, s3_generic
        ...
     (--) S3: card type: 386/486 localbus
     (--) S3: chipset:   864 rev. 0
     (--) S3: chipset driver: mmio_928

Here, we see that the two valid chipsets for this server (in this case, XF86 S3) are mmio 928 and s3 generic. The server probed for and found a video card using the mmio 928 chipset.

In the Device section of the XF86Config file, add a Chipset line, containing the name of the chipset as determined above. For example,

     Section "Device"
             # We already had Identifier here...
             Identifier "#9 GXE 64"
             # Add this line:
             Chipset "mmio_928"
     EndSection

Now we need to determine the driving clock frequencies used by the video card. A driving clock frequency, or dot clock, is simply a rate at which the video card can send pixels to the monitor. As we have seen, each monitor resolution has a dot clock associated with it. Now we need to determine which dot clocks are made available by the video card.

First you should look into the files (modeDB.txt, and so forth) mentioned above and see if your card’s clocks are listed there. The dot clocks will usually be a list of 8 or 16 values, all of which are in MHz. For example, when looking at modeDB.txt we see an entry for the Cardinal ET4000 video board, which looks like this:

     # chip    ram   virtual   clocks                           default-mode  flags
      ET4000   1024  1024 768   25  28  38  36  40  45  32   0  "1024x768"

As we can see, the dot clocks for this card are 25, 28, 38, 36, 40, 45, 32, and 0 MHz.

In the Devices section of the XF86Config file, you should add a Clocks line containing the list of dot clocks for your card. For example, for the clocks above, we would add the line

             Clocks 25 28 38 36 40 45 32 0

to the Devices section of the file, after Chipset. Note that the order of the clocks is important! Don’t resort the list of clocks or remove duplicates.

If you cannot find the dot clocks associated with your card, the X server can probe for these as well. Using the X -probeonly command described above, the output should contain lines which look like the following:

     (--) S3: clocks:  25.18  28.32  38.02  36.15  40.33  45.32  32.00  00.00

We could then add a Clocks line containing all of these values, as printed. You can use more than one Clocks line in XF86Config should all of the values (sometimes there are more than 8 clock values printed) not fit onto one line. Again, be sure to keep the list of clocks in order as they are printed.

Be sure that there is no Clocks line (or that it is commented out) in the Devices section of the file when using X -probeonly to probe for the clocks. If there is a Clocks line present, the server will not probe for the clocks—it will use the values given in XF86Config.

Note that some accelerated video boards use a programmable clock chip. (See the XF86 Accel man page for details; this generally applies to S3, AGX, and XGA-2 boards.) This chip essentially allows the X server to tell the card which dot clocks to use. If this is the case, then you may not find a list of dot clocks for the card in any of the above files. Or, the list of dot clocks printed when using X -probeonly will only contain one or two discrete clock values, with the rest being duplicates or zero.

For boards which use a programmable clock chip, you would use a ClockChip line, instead of a Clocks line, in your XF86Config file. ClockChip gives the name of the clock chip as used by the video card; the man pages for each server describe what these are. For example, in the file README.S3, we see that several S3-864 video cards use an “ICD2061A” clock chip, and that we should use the line

        ClockChip "icd2061a"

instead of Clocks in the XF86Config file. As with Clocks, this line should go in the Devices section, after

Chipset.

Similarly, some accelerated cards require you to specify the RAMDAC chip type in the XF86Config file, using a Ramdac line. The XF86 Accel man page describes this option. Usually, the X server will correctly probe for the RAMDAC.

Some video card types require you to specify several options in the Devices section of XF86Config. These options will be described in the man page for your server, as well as in the various files (such as README.cirrus or README.S3. These options are enabled using the Option line. For example, the #9 GXE 64 card requires two options:

             Option "number_nine"
             Option "dac_8_bit"

Usually, the X server will work without these options, but they are necessary to obtain the best performance. There are too many such options to list here, and they each depend on the particular video card being used. If you must use one of these options, fear not—the X server man pages and various files in /usr/X11R6/lib/X11/doc will tell you what they are.

So, when you’re finished, you should end up with a Devices section which looks something like this:

     Section "Device"
             # Device section for the #9 GXE 64 only!
             Identifier "#9 GXE 64"
             Chipset "mmio_928"
             ClockChip "icd2061a"
             Option "number_nine"
             Option "dac_8_bit"
     EndSection

Most video cards will require a Clocks line, instead of ClockChip, as described above. The above Device entry is only valid for a particular video card, the #9 GXE 64. It is given here only as an example.

There are other options that you can include in the Devices entry. Check the X server man pages for the gritty details, but the above should suffice for most systems.

6 Running XFree86

With your XF86Config file configured, you’re ready to fire up the X server and give it a spin. First, be sure that /usr/X11R6/bin is on your path.

The command to start up XFree86 is

     startx

This is a front-end to xinit (in case you’re used to using xinit on other UNIX systems).

This command will start the X server and run the commands found in the file .xinitrc in your home directory. .xinitrc is just a shell script containing X clients to run. If this file does not exist, the system

default /usr/X11R6/lib/X11/xinit/xinitrc will be used.

A standard .xinitrc file looks like this:

     #!/bin/sh
     xterm -fn 7x13bold -geometry 80x32+10+50 &
     xterm -fn 9x15bold -geometry 80x34+30-10 &
     oclock -geometry 70x70-7+7 &
     xsetroot -solid midnightblue &

     exec twm

This script will start up two xterm clients, an oclock, and set the root window (background) color to midnightblue. It will then start up twm, the window manager. Note that twm is executed with the shell’s exec statement; this causes the xinit process to be replaced with twm. Once the twm process exits, the X server will shut down. You can cause twm to exit by using the root menus: depress mouse button 1 on the desktop background—this will display a pop up menu which will allow you to Exit Twm.

Be sure that the last command in .xinitrc is started with exec, and that it is not placed into the background (no ampersand on the end of the line). Otherwise the X server will shut down as soon as it has started the clients in the .xinitrc file.

Alternately, you can exit X by pressing ctrl-alt-backspace in combination. This will kill the X server directly, exiting the window system.

The above is a very, very simple desktop configuration. Many wonderful programs and configurations are available with a bit of work on your .xinitrc file. For example, the fvwm window manager will provide a virtual desktop, and you can customize colors, fonts, window sizes and positions, and so forth to your heart’s content. Although the X Window System might appear to be simplistic at first, it is extremely powerful once you customize it for yourself.

If you are new to the X Window System environment, we strongly suggest picking up a book such as The X Window System: A User’s Guide. Using and configuring X is far too in-depth to cover here. See the man pages for xterm, oclock, and twm for clues on getting started.

7 Running Into Trouble

Often, something will not be quite right when you initially fire up the X server. This is almost always caused by a problem in your XF86Config file. Usually, the monitor timing values are off, or the video card dot clocks set incorrectly. If your display seems to roll, or the edges are fuzzy, this is a clear indication that the monitor timing values or dot clocks are wrong. Also be sure that you are correctly specifying your video card chipset, as well as other options for the Device section of XF86Config. Be absolutely certain that you are using the right X server and that /usr/X11R6/bin/X is a symbolic link to this server.

If all else fails, try to start X “bare”; that is, use a command such as:

     X > /tmp/x.out 2>&1

You can then kill the X server (using the ctrl-alt-backspace key combination) and examine the contents of /tmp/x.out. The X server will report any warnings or errors—for example, if your video card doesn’t have a dot clock corresponding to a mode supported by your monitor.

The file VideoModes.doc included in the XFree86 distribution contains many hints for tweaking the values

in your XF86Config file.

Remember that you can use ctrl-alt-numeric + and ctrl-alt-numeric - to switch between the video modes listed on the Modes line of the Screen section of XF86Config. If the highest resolution mode doesn’t look right, try switching to lower resolutions. This will let you know, at least, that those parts of your X configuration are working correctly.

Also, check the vertical and horizontal size/hold knobs on your monitor. In many cases it is necessary to adjust these when starting up X. For example, if the display seems to be shifted slightly to one side, you can usually correct this using the monitor controls.

The USENET newsgroup comp.windows.x.i386unix is devoted to discussions about XFree86, as is comp.os.linux.x. It might be a good idea to watch that newsgroup for postings relating to your video configuration—you might run across someone with the same problems as your own.

8 Copyright

This document is Copyright (c)1995 by Matt Welsh. This work may be reproduced and distributed in whole or in part, in either printed or electronic form, subject to the following conditions:

1. The copyright notice and this license notice must be preserved complete on all complete or partial copies.

2. Any translation or derivative work must be approved by the author in writing before distribution.

3. If you distribute the Work in part, instructions for obtaining a complete version (in printed or electonic form) must be included, and a means for obtaining a complete version provided.

4. Small portions may be reproduced as illustrations for reviews or quotes in other works without this permission notice if proper citation is given.

Exceptions to these rules may be granted for academic purposes, write to the author of the Work, and ask. These restrictions are here to protect the authors, not to restrict you as educators and learners.